.1.3.6.1.2.1.2.2.1
ifIndex (1)_{1   2   3   4   5}

ifDescr (2)_{1   2   3   4   5}

ifType (3)_{1   2   3   4   5}

ifMtu (4)_{1   2   3   4   5}

ifSpeed (5)_{1   2   3   4   5}

ifPhysAddress (6)   ifAdminStatus (7)   ifOperStatus (8)   ifLastChange (9)    ifInOctets (10)   ifInUcastPkts (11)   ifInNUcastPkts (12)   ifInDiscards (13)   ifInErrors (14)   ifInUnknownProtos (15)   ifOutOctets (16)   ifOutUcastPkts (17)   ifOutNUcastPkts (18)   ifOutDiscards (19)   ifOutErrors (20)    ifOutQLen (21)     ifSpecific (22)
1   2   3   4   5   1   2   3   4   5   1   2   3   4   5   1   2   3   4   5   1   2   3   4   5   1   2   3   4   5     1   2   3   4   5    1   2   3   4   5   1   2   3   4   5     1   2   3   4   5     1   2   3   4   5    1   2   3   4   5      1   2   3   4   5     1   2   3   4   5   1   2   3   4   5   1   2   3   4   5   1   2   3   4   5